) Six years, ago, the ratio of the ages of kamala and Suresh was 6:5. Four years hence the ratio of their ages will be 11:10. What is Suresh’s age at present?
Answer by mananth(16946)

Suresh age now = x
Kamala age now = y
6 years ago
suresh age = x-6
Kamala age = y-6
(y-6)/(x-6)=6/5
5(y-6)=6(x-6)
5y-30=6x-36
5y-6x=-6..............1
after 4 years
suresh x+4
Kamala y+4
(y+4)/(x+4)=11/10
10(y+4)=11(x+4)
10y+40=11x+44
10y-11x=4.............2

-6 x + 5 y = -6 .............1
-11 x + 10 y = 4 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
12 x -10 y = 12
-11 x 10 y = 4
Add the two equations
x = 16
plug value of x in (1)
-6 x + 5 y = -6
-96 + 5 y = -6
5 y = -6 + 96
5 y = 90
y = 18
suresh's age = 16 years